OpenStudy (anonymous):

Is (-1, -1) a solution of y = 3x + 2?

OpenStudy (precal):

-1=3(-1)+2 -1=-3+2 -1=-1 true

OpenStudy (saifoo.khan):

Try inserting the values, -1 = 3(-1) + 2 THe left side should be equal to right side.